OpenStudy (ladiesman217):

medals & fans rewarded Create a quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (ladiesman217):

@.Sam. can you help me please

OpenStudy (anonymous):

Why is the value of arcsin(sin(5pi/4)) -pi/4 and not 5pi/4? Thank you. :)

OpenStudy (compassionate):

-16t^2 + 8(t) + 15 (x - 3)(x - 5) Points (3, 5) graphing. Since 3 is directly correlated to five, plug in the numbers and you'l get. 

OpenStudy (compassionate):

-16t^2 + 8(t) + 15 (x - 3)(x - 5) Points (3, 5) graphing. The binomial terms are the roots. I can then put it in a function y = -16t^2 + 8(t) + 15, substituting -3, 3, which will give me the parabola. To use the quadratic equation: \[\frac{ - 8 \pm \sqrt{8^2 - 4(-16)(15)}}{ 2(-16) } \] Solving the two roots, will give me 3,5. I can also use this formula to maintain the minimum or maximum by using the equation. \[\frac{ -b }{ 2a }\]